Suppose Cho is a cinephile and buys only movie tickets. Cho deposits $2,000 in a bank account that pays an annual nominal interest rate of 15% Assume this interest rate is fixed-that is, it wont change over time. At the time of her deposit, a movie ticket is priced at $20.00 Initially, the purchasing power of Chos $2,000 deposit ismovie tickets. For each of the annual inflation rates given in the following table, first determine the new price of a movie ticket, assuming it rises at the rate of inflation. Then enter the corresponding purchasing power of Chos deposit after one year in the first row of the table for each inflation rate. Finally, enter the value for the real interest rate at each of the given inflation rates. Hint: Round your answers in the first row down to the nearest movie ticket. For example, if you find that the deposit will cover 20.7 movie tickets, you would round the purchasing power down to 20 movie tickets under the assumption that Cho will not buy seven-tenths of a movie ticket.
Answer #1

Initially, the purchasing power of Cho's $2000 deposit = $2000 / $20 = 100 movie tickets

Annual inflation rate
0% 15% 18%
Number of tickets Cho can purchase after one year 115 100 97
Real interest rate 15% 0% -3%

Explanation:

Real interest rate = Nominal interest rate - Inflation rate

  • The value of $2000 after one year = 2000(1 + 0.15) = $2300. Since inflation rate is 0%, the price will remain the same.So, the Number of tickets Cho can purchase after one year = 2300 / 20 = 115 movie tickets
  • The value of $2000 after one year = 2000(1 + 0.15) = $2300. Since inflation rate is 15%, the price of the ticket = 20 + (20 * 0.15) = $23. So, the Number of tickets Cho can purchase after one year = 2300 / 23 = 100 movie tickets.
  • The value of $2000 after one year = 2000(1 + 0.15) = $2300. Since inflation rate is 18%, the price of the ticket = 20 + (20 * 0.18) = $23.60. So, the Number of tickets Cho can purchase after one year = 2300 / 23.60 = 97 movie tickets.

Answer to the blank: remains the same